what is an equation of the perpendicular bisector of the line between the points, (2,2) and (6,6). im not sure how to get this. please help me if possible thank you
Answer by Alan3354(69443)

what is an equation of the perpendicular bisector of the line between the points, (2,2) and (6,6).
---------------
The slope of the line thru the points is 1 (6-2)/(6-2)
The midpoint is (4,4)
---------------
The slope of the line perpendicular is -1
Use y = mx+b to find the line thru (4,4) and m = -1
4 = -4 + b
b = 8
------
--> y = -x + 8
or x + y = 8